# Ultra-Fast Language Generation via Discrete Diffusion Divergence Instruct

### v1

@@ 4.1 Discrete Diffusion Divergence Instruction
where the expectation over time $t$ is sampled from distribution $\pi(t)$, and $R(\mathbf{z}_{t},t):=\log\mathbf{p}_{\nu}(\mathbf{z}_{t},t)-\log\mathbf{p}_{\theta}(\mathbf{z}_{t},t)$ denotes the reward (log-density ratio between student and teacher) at $\mathbf{z}_{t}$.
@@ 4.2 Density‑Ratio Estimation
where $D_{\lambda}^{\ell}(\cdot,\cdot)$ denotes the $\ell$-th element of $D_{\lambda}(\cdot,\cdot)$. This procedure provides a tractable and computable reward signal for our objective according to $D_{\lambda}$.
@@ 4.4 The Proposed Algorithms
Figure 2: The pipeline of DiDi-Instruct (Algorithm 1). Given a fully masked input $\mathbf{z}_{t}$ ($t=1$), both the student $\mathbf{p}_{\nu}$ and the teacher (assistant model) $\mathbf{p}_{\theta}$ produce clean samples $\mathbf{x}$ and $\mathbf{x}^{\prime}$, which are corrupted at $t_{i}\sim\pi(t)$ to form $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{\prime}$. The discriminator $D_{\lambda}$ is trained to classify these outputs, while its reward signal (6) enables the gradient update (5) for the student. The red line represents the gradient flow for the student’s update step, and the blue line represents the gradient flow for the auxiliary model’s update step.
End-to-end training algorithm. The DiDi-Instruct training procedure alternates between updating the discriminator $D_{\lambda}$ and the student $\mathbf{p}_{\nu}$, which is summarized in Algorithm 1 and visualized in Figure 2. Each iteration consists of two phases: first, the discriminator is updated to better distinguish between corrupted samples from $\mathbf{p}_{\theta}$ and $\mathbf{p}_{\nu}$; we apply (5) with the normalized reward and the decomposed score to stabilize gradient updates. This alternating scheme yields a robust few-step generator that closely matches the teacher’s marginals over different corrupt levels.
@@
Reward-guided ancestral sampler. We further propose a decoding strategy that leverages the trained discriminator to guide ancestral sampling (AS) (Shi et al., 2024; Zheng et al., 2025). Starting from a fully masked sequence $\mathbf{z}_{N}=(\mathbf{m},\dots,\mathbf{m})$ at $t_{N}=1$, the procedure generates samples by iteratively denoising from $t_{n}$ to $t_{n-1}$ for $n=N,\dots,1$, following the student’s backward distribution $\mathbf{p}_{\nu}(\mathbf{z}_{n-1}|\mathbf{z}_{n})$. Specifically, for $\mathbf{z}_{n}^{\ell}\in\mathcal{V}$ ($\ell=1,\cdots,L$) we consider a tilted transition:
@@ 5 Experiments
Our experiments are designed to distill a pre-trained teacher model into an efficient few-step student generator with DiDi-Instruct. All models are trained on OWT (Gokaslan et al., 2019). Following standard practices (Sahoo et al., 2025), we tokenize the corpus using the GPT-2 tokenizer, pack sequences to a context length of 1024, and hold out the last 100,000 documents for validation.
Our teacher is a 169M parameter MDLM with a Diffusion Transformer (Peebles and Xie, 2023) (12 layers, 12 attention heads, 768 hidden dimension). We pre-trained this model from scratch for 1024 NFEs, achieving a perplexity of 38.53 and an entropy of 5.22. The student model shares an identical architecture to ensure a fair comparison. The reward model is a 131M parameter network based on the same backbone, but with a new randomly initialized classification head. This head consists of two linear layers with spectral normalization and a SiLU activation function. During distillation, the reward model and the student generator are trained adversarially in an alternating fashion.
@@ 5.1 Experimental Result and Analysis
Generation results. We evaluate generative quality by sampling text from the distilled student from $8$ to $128$ NFEs and measuring GPT-2 Large generative PPL and average sequence entropy. Figure 1 shows that DiDi-Instruct consistently outperforms all baselines in PPL across all sampling steps. Notably, with only 16 NFEs, our model’s PPL already surpasses that of the 1024-step teacher model. At 1024 NFEs, DiDi-Instruct achieves a final PPL of 15.62, a reduction of over 24% compared to the strongest baseline. These performance gains are achieved with a negligible loss in diversity; the generative entropy from 8 to 128 NFEs is 5.17, 5.21, 5.18, 5.15, and 5.15, respectively (e.g., 1024-step teacher model is 5.22), indicating sample diversity is well-preserved.
Efficiency-performance tradeoff. DiDi-Instruct offers substantial computational advantages in both training and inference. Our single-round distillation framework completes training in around one H100 GPU hour, in contrast to the 2020+ GPU hours required by multi-round methods (Sahoo et al., 2024; Deschenaux and Gulcehre, 2025). During inference, RGAS provides a superior perplexity-latency trade-off. As shown in Figure 3, it achieves a PPL of $23.54$ at a latency of $1.31$ s/sequence ($32$ NFEs), whereas the ancestral sampler only reaches a PPL of $44.54$ at a similar latency ($0.71$ s/sequence, $32$ NFEs).
@@
Sample quality. Appendix F.7 presents full text excerpts with per-sample metrics (PPL and entropy) for the 1024-step teacher and DiDi-Instruct students from 8 to 128 NFEs, which confirms a clear improvement in narrative quality. The 8-step student exhibits expected repetition, a common artifact of rapid sampling. However, this issue is resolved by 16 NFEs, at which point the student model already surpasses the 1024-step teacher in paragraph-level coherence and topic adherence. This ability to construct focused and specific narratives strengthens as NFEs increase to 128, indicating that DiDi-Instruct not only preserves fluency but actively enhances the model’s ability to generate structured, coherent text.
@@ 5.3 Model Scaling Up
We perform an incremental scaling experiment to 424M parameters while holding the training and inference pipeline fixed to the 169M configuration. The teacher model is a pre-trained 424M MDLM, which is a DiT-style transformer with a hidden size of 1024, 24 blocks, and 16 attention heads (check Table 3 for detailed network structure). The student is distilled on a single H100 using the same data, masking policy, optimizer, and schedule. To balance the training, we also scale the reward discriminator to 373M and employ a deeper network with more trainable parameters.
The generative PPL and entropy are shown in Figure 4, and we report improvements relative to the teacher at matched NFEs here: from 8 to 1024 NFEs, DiDi-Instruct yields large and consistent PPL reductions of $88.5\%$ (8 NFEs), $87.7\%$ (16 NFEs), $85.3\%$ (32 NFEs), $82.7\%$ (64 NFEs), and $79.0\%$ (128 NFEs). Moreover, with only 16 NFEs, the distilled model achieves a perplexity of 32.79, marking an 11.4% improvement over the 1024-step teacher baseline. Entropy remains comparable throughout, indicating that diversity is preserved while accuracy improves. These results confirm that the quality–efficiency advantages of DiDi-Instruct persist at increasing scale with minimal procedural changes. For more details, please refer to Appendix F.5.
@@ 6 Conclusion and Future Work
In this work, we introduced DiDi-Instruct, a training-based acceleration framework for fast language generation that distills a high-quality teacher into a few-step student. Our design targets three axes simultaneously: (i) objective design via a tractable policy–gradient update driven by a discriminator–estimated reward; (ii) training stability through score decomposition and grouped reward normalization; and (iii) inference efficiency through reward-guided ancestral sampling with gradient tilting and re-ranking. Experiments demonstrate strong gains in generation quality, large reductions in training/inference time, and competitive zero-shot generalization, corroborated by comprehensive cumulative and leave-one-out ablations.
We plan to scale DiDi-Instruct to billion-parameter models, which presents a practical challenge due to the memory requirements of concurrently maintaining the teacher, student, and discriminator. Nevertheless, our findings already establish a new state-of-the-art trade-off among comparable methods, with the student model excelling in quality at low NFEs and maintaining diversity at higher computational budgets. We posit that DiDi-Instruct offers a foundational recipe (principled objectives, training stability, and efficient guidance) for developing high-performance generative models.
@@ Appendix A Language Generation Showcase
To provide a qualitative illustration of our method’s effectiveness, we present a series of uncurated Samples from model below. The first sample is generated from the 1024-step teacher model, serving as a high-quality reference. Subsequent samples are generated by our distilled model, DiDi-Instruct, with an increasing NFEs from 8 to 128. These examples clearly demonstrate the progressive improvement in coherence and semantic quality as the sampling budget increases, showcasing how DiDi-Instruct efficiently approaches the teacher’s performance. For more examples, please refer to Appendix F.7.
Samples from model obtained with $1024$ NFEs teacher model. Preplexity=$40.48$, Entropy=$5.22$. He parked and it was then that I saw that president and ended up there, speaking with a journalist. The Prince was granted anonymity and noted that the FBI was aware of the familiarity with Russia and was conducting the investigation as appropriate. Victor Atkinson, an FBI chief-officer who is retired, hacked the personal PINs and cell phone lines of the two Trump officials to conceal the information, as shown inside the countrys electronic networks. Anonymous messages can appear in our apps like Twitter and Facebook.
@@ B.1 Related Work on Continuous-time Diffusion Models
Continuous-space diffusion distillation. Diffusion models (Song et al., 2021b; Ho et al., 2020; Sohl-Dickstein et al., 2015) added Gaussian noises to data with a forward diffusion process and then learn the backward stochastic differential equation (SDE) with so-called score neural networks. Early acceleration reduced the steps of iterative samplers using training-free high-order ODE solvers (Song et al., 2021a; Lu et al., 2022; Karras et al., 2022; Xue et al., 2023). However, solver-based acceleration suffered from discretization error in less than 10 generation steps.
@@ B.3 Related Work on Few-Step Discrete Diffusion LLMs
To distill discrete large language models (dLLMs) into few-step generators for fast language generation, a natural approach is to extend the well-developed distillation techniques already established for continuous diffusion models to the discrete domain. However, in MDMs, all tokens are eventually mapped to the masked state, meaning the prior distribution collapses to the fully masked sequence. As a result, generation under MDMs is inherently stochastic, in contrast to the deterministic trajectories available in the continuous case. Consequently, the stochasticity in Masked Diffusion Models (MDMs) arises from the choice of the masking schedule; once a schedule is fixed, the reverse path is deterministic. This stands in sharp contrast to the inherently stochastic reverse dynamics of continuous diffusion. This determinism is problematic for distillation, as training a student model solely on these fixed trajectories risks mode collapse and a failure to capture the teacher’s generalization capabilities. This fundamental difference makes the extensive body of work on continuous-state diffusion distillation not directly applicable. To make this distinction concrete, we first show how the following methods address these challenges in discrete diffusion distillation, followed by a comparison of these representative approaches.
SDTT (Deschenaux and Gulcehre, 2025).
@@
SDTT progressively distills a multistep teacher diffusion into a fewer-step student model. Due to the absence of an ODE trajectory as in continuous cases, it matches the teacher-student distribution via KL divergence minimization. Similar to DUO (Sahoo et al., 2025), it needs to decrease the number of student steps for training stability gradually. Such a curriculum-based strategy can be inefficient and computationally costly.
DUO (Sahoo et al., 2025).
@@
Continuous-space consistency distillation (CD) requires probability flow ODE (PF-ODE) trajectories, while MDMs are inherently stochastic, making direct CD inapplicable. DUO overcomes this by connecting Gaussian diffusion with Uniform-state Diffusion Model (USDM), enabling CD in the discrete setting. By applying an $\operatorname*{arg\,max}$ to the latent vectors of Gaussian diffusion, continuous vectors are mapped into discrete one-hot tokens. It is proven that the resulting marginal distribution exactly matches that of a USDM under a suitably transformed noise schedule. Consequently, the evolution of a USDM can be described by an ODE, which makes CD feasible. However, performing CD on USDM requires multiple rounds of distillation and a carefully annealed discretization schedule for training stability. Specifically, the CD procedure starts from the USDM diffusion and gradually learns to map two points, $t$ and $t-\Delta t$, along the same ODE trajectory back to the same origin. Initially, since the diffusion model cannot take large steps without losing accuracy, $\Delta t$ must be kept small. Once the model learns to consistently align $t$ and $t-\Delta t$, the step size can be increased, eventually reaching the largest possible step $\Delta t=t$, i.e., the distilled model can jump large steps for few-step inference. A small step is safe but biased, while a large step is unbiased but unstable. This tradeoff is similar to the observation of CD in the continuous domain.
DiMO (Zhu et al., 2025b).
@@
DiMO distills multi-step MDMs into a one-step generator. The key idea is to augment the prior distribution: instead of restricting the initial state to a fully-masked sequence, DiMO samples a subset of tokens from the entire vocabulary. This relaxation enriches the prior, allowing the model to leverage partial information during generation and improving both efficiency and diversity of outputs. During training, DiMO employs an on-policy distillation strategy. The one-step generator is supervised to match the teacher’s conditional prediction distribution, not its final output distribution. Specifically, it uses the generator’s own one-step output to create a pseudo-intermediate state, and then minimizes the divergence between the student’s and the teacher’s predicted token distributions conditioned on this state. To avoid mode collapse and reduce mismatch with the teacher’s training distribution, DiMO introduces a token initialization strategy that mixes mask tokens with random tokens and adds Gaussian perturbations to embeddings.
@@
Our advantages over three related methods.
Our work presents a straightforward and intuitive framework for distilling dLLMs by directly matching teacher and student distributions, addressing several key limitations of recent approaches. Unlike multi-stage progressive distillation techniques (Sahoo et al., 2025; Deschenaux and Gulcehre, 2025), our method is a direct, single-stage process. It not only avoids the use of deterministic or heuristic mappings, but also dispenses with the compute-intensive, multi-round training used by consistency-based approaches, which requires careful choice of time discretization step annealing. Furthermore, by aligning the student with the teacher’s entire stochastic process, our framework offers broad applicability to general dLLMs, in contrast to methods tailored for specific architectures, e.g., DUO (Sahoo et al., 2025) relies on USDM for duality. Finally, we provide a mathematically rigorous solution to the problem of non-differentiability in discrete spaces. This avoids the need for biased proxy-gradient estimators, whose performance in purely textual domains remains underexplored (Zhu et al., 2025b), and instead offers a principled path for distillation.

### v2

@@ 4.1 Discrete Diffusion Divergence Instruction
where the expectation over time $t$ is sampled from distribution $\pi(t)$, and $R(\mathbf{z}_{t},t):=\log\mathbf{q}_{\nu}(\mathbf{z}_{t},t)-\log\mathbf{q}_{\theta}(\mathbf{z}_{t},t)$ denotes the reward (log-density ratio between student and teacher) at $\mathbf{z}_{t}$.
@@ 4.2 Density‑Ratio Estimation
where $D_{\lambda}^{\ell}(\cdot,\cdot)$ denotes the $\ell$-th element of $D_{\lambda}(\cdot,\cdot)$ ($\ell=1,2,\cdots,L$). This procedure provides a tractable and computable reward signal for our objective according to $D_{\lambda}$.
@@ 4.4 The Proposed Algorithms
Figure 2: The pipeline of DiDi-Instruct (Algorithm 1). Given a fully masked input $\mathbf{z}_{t}$ ($t=1$), both the student $\mathbf{p}_{\nu}$ and the teacher $\mathbf{p}_{\theta}$ produce clean samples $\mathbf{x}$ and $\mathbf{x}^{\prime}$, which are corrupted at $t_{i}\sim\pi(t)$ to form $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{\prime}$. The discriminator $D_{\lambda}$ is trained to classify these outputs, while its reward signal (6) enables the gradient update (5) for the student. The red line denotes the gradient flow for the student’s update step, and the blue line represents the one for the auxiliary model’s update step.
End-to-end training algorithm. The DiDi-Instruct training procedure alternates between updating the discriminator $D_{\lambda}$ and the student $\mathbf{p}_{\nu}$, which is summarized in Algorithm 1 and visualized in Figure 2. Each iteration consists of two phases: first, the discriminator is updated to better distinguish between corrupted samples between $\mathbf{p}_{\theta}$ and $\mathbf{p}_{\nu}$; we apply (5) with the normalized reward and the decomposed score to stabilize gradient updates. This alternating scheme yields a robust few-step generator that closely matches the teacher’s marginals over different corrupt levels.
@@
Reward-guided ancestral sampler. We further propose a decoding strategy that leverages the trained discriminator to guide ancestral sampling (AS) (Shi et al., 2024; Zheng et al., 2025). Starting from a fully masked sequence $\mathbf{z}_{N}=(\mathbf{m},\dots,\mathbf{m})$ at $t_{N}=1$, the procedure generates samples by iteratively denoising from $t_{n}$ to $t_{n-1}$ for $n=N,\dots,1$, following the student’s backward distribution $\mathbf{p}_{\nu}(\mathbf{z}_{n-1}|\mathbf{z}_{n})$. Specifically, for $\mathbf{z}_{n}^{\ell}\in\mathcal{V}$ ($\ell=1,\cdots,L$), we consider a tilted transition:
@@ 5 Experiments
Our experiments are designed to distill a pre-trained teacher model into an efficient few-step student generator with DiDi-Instruct. All models are trained on OpenWebText (Gokaslan et al., 2019). Following standard practices (Sahoo et al., 2025), we tokenize the corpus using the GPT-2 tokenizer, pack sequences to a context length of 1024, and hold out the last 100,000 documents for validation.
Our teacher is a 169M parameter MDLM with a Diffusion Transformer (Peebles and Xie, 2023) (12 layers, 12 attention heads, 768 hidden dimension). We pre-trained this model for 1024 NFEs, achieving a perplexity of 38.53 and an entropy of 5.22. The student model shares an identical architecture to ensure a fair comparison. The discriminator is a 131M parameter network based on the same backbone, but with a new randomly initialized classification head. This head consists of two linear layers with spectral normalization and a SiLU activation function. During distillation, the discriminator model and the student generator are trained adversarially in an alternating fashion.
@@ 5.1 Experimental Result and Analysis
Generation results. We evaluate generative quality by sampling text from the distilled student from $8$ to $128$ NFEs and measuring GPT-2 Large generative PPL and average sequence entropy. Figure 1 shows that DiDi-Instruct consistently outperforms all baselines in PPL across all sampling steps. Notably, with only 16 NFEs, our model’s PPL already surpasses that of the 1024-step teacher model. At 1024 NFEs, DiDi-Instruct achieves a final PPL of 15.62, a reduction of over 24% compared to the strongest baseline. These performance gains are achieved with a negligible loss in diversity; the generative entropy from 8 to 128 NFEs is 5.17, 5.22, 5.20, 5.18, and 5.17, respectively (e.g., 1024-step teacher model is 5.22), indicating sample diversity is well-preserved.
Efficiency-performance tradeoff. DiDi-Instruct offers substantial computational advantages in both training and inference. Our single-round distillation framework completes training in around one H100 GPU hour, in contrast to the 2020+ GPU hours required by multi-round methods (Sahoo et al., 2024; Deschenaux and Gulcehre, 2025). During inference, RGAS provides a superior perplexity-latency trade-off. As shown in Figure 3, it achieves a PPL of $23.54$ at a latency of $1.31$ s/sequence ($32$ NFEs), whereas AS only reaches a PPL of $44.54$ ($0.71$ s/sequence, $32$ NFEs).
@@
Sample quality. Appendix F.7 presents full text excerpts with per-sample metrics (PPL and entropy) for the 1024-step teacher and DiDi-Instruct students from 8 to 128 NFEs, which confirms a clear improvement in narrative quality. The 8-step student exhibits expected repetition, a common artifact of rapid sampling. However, this issue is resolved by 16 NFEs, at which point the student model already surpasses the 1024-step teacher in paragraph-level coherence and topic adherence. This ability to construct focused and specific narratives strengthens as NFEs increase to 128, indicating that DiDi-Instruct not only preserves fluency, but actively enhances the model’s ability to generate structured and coherent text.
@@ 5.3 Model Scaling Up
We perform an incremental scaling experiment to 424M parameters while holding the training and inference pipeline fixed to the 169M configuration. The teacher model is a pre-trained 424M MDLM, which is a Diffusion Transformer with a hidden size of 1024, 24 blocks, and 16 attention heads (check Table 4 for detailed network structure). The student is distilled on a single H100 using the same data, masking policy, optimizer, and schedule. To balance the training, we also scale the discriminator to 373M and employ a deeper network with more trainable parameters.
The generative PPL and entropy are shown in Figure 4, and we report improvements relative to the teacher at matched NFEs here: from 8 to 1024 NFEs, DiDi-Instruct yields large and consistent PPL reductions of $88.5\%$ (8 NFEs), $87.7\%$ (16 NFEs), $85.3\%$ (32 NFEs), $82.7\%$ (64 NFEs), and $79.0\%$ (128 NFEs). Entropy remains comparable throughout, indicating that diversity is preserved while accuracy improves. These results confirm that the quality–efficiency advantages of DiDi-Instruct persist at increasing scale with minimal procedural changes. For more details, please refer to Appendix F.5.
@@ 6 Conclusion and Future Work
In this work, we introduced DiDi-Instruct, a training-based acceleration framework for fast language generation that distills a high-quality teacher into a few-step student. Our design targets three axes simultaneously: (i) objective design via a tractable policy–gradient update driven by a discriminator–estimated reward; (ii) training stability through score decomposition and grouped reward normalization; and (iii) inference efficiency through reward-guided ancestral sampling with gradient tilting and re-ranking. Experiments demonstrate strong gains in generation quality, large reductions in training/inference time, and competitive zero-shot generalization, corroborated by comprehensive cumulative, leave-one-out ablations, model scaling up, and protein sequence generation.
We plan to scale DiDi-Instruct to billion-parameter models, which presents a practical challenge due to the GPU memory requirements of concurrently maintaining the teacher, student, and discriminator. Nevertheless, our findings already establish a new state-of-the-art trade-off among comparable methods, with the student model excelling in quality at low NFEs and maintaining diversity at higher computational budgets. We posit that DiDi-Instruct offers a foundational recipe (principled objectives, training stability, and efficient guidance) for developing high-performance generative models.
@@ Appendix A Language Generation Showcase
To provide a qualitative illustration of our method’s effectiveness, we present a series of uncurated samples from the base and distilled model below. The first sample is generated from the 1024-step teacher model, serving as a high-quality reference. Subsequent samples are generated by DiDi-Instruct, with an increasing NFEs from 8 to 128. These examples clearly demonstrate the progressive improvement in coherence and semantic quality as the sampling budget increases, showcasing how DiDi-Instruct efficiently approaches the teacher’s performance. For more examples, please refer to Appendix F.7.
Samples from model obtained with $1024$ NFEs teacher model. Preplexity=$38.53$, Entropy=$5.22$. He parked and it was then that I saw that president and ended up there, speaking with a journalist. The Prince was granted anonymity and noted that the FBI was aware of the familiarity with Russia and was conducting the investigation as appropriate. Victor Atkinson, an FBI chief-officer who is retired, hacked the personal PINs and cell phone lines of the two Trump officials to conceal the information, as shown inside the countrys electronic networks. Anonymous messages can appear in our apps like Twitter and Facebook.
@@ B.1 Related Work on Continuous-Time Diffusion Models
Diffusion models (Song et al., 2021b; Ho et al., 2020; Sohl-Dickstein et al., 2015) added Gaussian noises to data with a forward diffusion process and then learn the backward stochastic differential equation (SDE) with so-called score neural networks. Early acceleration reduced the steps of iterative samplers using training-free high-order ODE solvers (Song et al., 2021a; Lu et al., 2022; Karras et al., 2022; Xue et al., 2023). However, solver-based acceleration suffered from discretization error in less than 10 generation steps.
@@ B.3 Related Work on Few-Step Discrete Diffusion LLMs
To distill dLLMs into few-step generators for fast language generation, a natural approach is to extend the well-developed distillation techniques already established for continuous diffusion models to the discrete domain. However, in MDMs, all tokens are eventually mapped to the masked state, meaning the prior distribution collapses to the fully masked sequence. As a result, generation under MDMs is inherently deterministic, in contrast to the stochastic reverse dynamics in continuous diffusion. Consequently, the stochasticity in MDMs arises from the choice of the masking schedule; once a schedule is fixed, the reverse path is deterministic. This stands in sharp contrast to the inherently stochastic reverse dynamics of continuous diffusion. This determinism is problematic for distillation, as training a student model solely on these fixed trajectories risks mode collapse and a failure to capture the teacher’s generalization capabilities. This fundamental difference makes the extensive body of work on continuous-state diffusion distillation not directly applicable. To make this distinction concrete, we first show how the following methods address these challenges in discrete diffusion distillation, followed by a comparison of these representative approaches.
@@
SDTT (Deschenaux and Gulcehre, 2025). SDTT progressively distills a multistep teacher diffusion into a fewer-step student model. Due to the absence of an ODE trajectory as in continuous cases, it matches the teacher-student distribution via KL divergence minimization. Similar to DUO (Sahoo et al., 2025), it needs to decrease the number of student steps for training stability gradually. Such a curriculum-based strategy can be inefficient and computationally costly.
@@
DUO (Sahoo et al., 2025). Continuous-space consistency distillation (CD) requires probability flow ODE (PF-ODE) trajectories, while MDMs are inherently stochastic, making direct CD inapplicable. DUO overcomes this by connecting Gaussian diffusion with USDMs, enabling CD in the discrete setting. By applying an $\operatorname*{arg\,max}$ to the latent vectors of Gaussian diffusion, continuous vectors are mapped into discrete one-hot tokens. It is proven that the resulting marginal distribution exactly matches that of a USDM under a suitably transformed noise schedule. Consequently, the evolution of a USDM can be described by an ODE, which makes CD feasible. However, performing CD on USDM requires multiple rounds of distillation and a carefully annealed discretization schedule for training stability. Specifically, the CD procedure starts from the USDM diffusion and gradually learns to map two points, $t$ and $t-\Delta t$, along the same ODE trajectory back to the same origin. Initially, since the diffusion model cannot take large steps without losing accuracy, $\Delta t$ must be kept small. Once the model learns to consistently align $t$ and $t-\Delta t$, the step size can be increased, eventually reaching the largest possible step $\Delta t=t$, i.e., the distilled model can jump large steps for few-step inference. A small step is safe but biased, while a large step is unbiased but unstable. This tradeoff is similar to the observation of CD in the continuous domain.
@@
DiMO (Zhu et al., 2025b). DiMO distills multi-step MDMs into a one-step generator. The key idea is to augment the prior distribution: instead of restricting the initial state to a fully-masked sequence, DiMO samples a subset of tokens from the entire vocabulary. This relaxation enriches the prior, allowing the model to leverage partial information during generation and improving both efficiency and diversity of outputs. During training, DiMO employs an on-policy distillation strategy. The one-step generator is supervised to match the teacher’s conditional prediction distribution, not its final output distribution. Specifically, it uses the generator’s own one-step output to create a pseudo-intermediate state, and then minimizes the divergence between the student’s and the teacher’s predicted token distributions conditioned on this state. To avoid mode collapse and reduce mismatch with the teacher’s training distribution, DiMO introduces a token initialization strategy that mixes mask tokens with random tokens and adds Gaussian perturbations to embeddings.
@@
Advantages of DiDi-Instruct. Our work presents a straightforward and intuitive framework for distilling dLLMs by directly matching teacher and student distributions, addressing several key limitations of recent approaches. Unlike multi-stage progressive distillation techniques (Sahoo et al., 2025; Deschenaux and Gulcehre, 2025), our method is a direct, single-stage process. It not only avoids the use of deterministic or heuristic mappings, but also dispenses with the compute-intensive, multi-round training used by consistency-based approaches, which requires careful choice of time discretization step annealing. Furthermore, by aligning the student with the teacher’s entire stochastic process, our framework offers broad applicability to general dLLMs, in contrast to methods tailored for specific architectures, e.g., DUO (Sahoo et al., 2025) relies on USDM for duality. Finally, we provide a mathematically rigorous solution to the problem of non-differentiability in discrete spaces. This avoids the need for biased proxy-gradient estimators, whose performance in purely textual domains remains underexplored (Zhu et al., 2025b), and instead offers a principled path for distillation.